# A Comprehensive Survey of Website Fingerprinting Attacks and Defenses in Tor: Advances and Open Challenges

### v1

@@ V-B Survey on WF Defenses
Tor encrypts and routes traffic through a network of relays to obscure user identity and destination, but it does not conceal traffic bursts or timing information, leaving it vulnerable to WF attacks. These attacks are particularly concerning in Tor, where users rely on strong privacy guarantees [159]. WF attacks against the first-hop (entry) relay can accurately infer user activity, especially in closed-world settings where the set of possible websites is predefined. To address this threat, researchers have proposed various WF defenses aimed at reducing the distinguishability of traffic patterns across websites, thereby diminishing the effectiveness of WF classifiers.
@@ V-B1 Adaptive Padding-Based Defenses
Adaptive traffic padding is a technique used to protect against WF attacks by dynamically injecting dummy traffic and introducing controlled delays into network traffic streams [53]. This dynamic behavior can be achieved by traffic obfuscation or randomized padding. Traffic obfuscation involves inserting dummy packets into statistically improbable timing gaps or low-activity regions of the traffic stream, making it more difficult for an attacker to identify meaningful patterns. Randomized padding further enhances this effect by varying the timing and size of the inserted packets, thereby increasing uncertainty for WF classifiers. This adaptive approach enables the system to disrupt fingerprinting attacks while maintaining low overhead in terms of bandwidth and latency [77]. Figure 7 illustrates the mechanism of adaptive padding. In the figure, coral rectangles represent outgoing packets, while light blue rectangles denote incoming packets. The light yellow and light green rectangles indicate the dummy packets added for outgoing and incoming traffic, respectively. The insertion of these dummy packets helps obscure burst patterns, timing information, and other features commonly exploited by WF attacks.
@@
TABLE X summarizes existing representative adaptive padding-based defenses, highlighting their strengths and limitations to facilitate comparison for researchers. Juarez et al. [16] introduced WTF-PAD, an adaptive padding defense that protects Tor users from WF attacks. This defense employs adaptive padding to disrupt distinctive traffic patterns, thereby significantly reducing the accuracy of traditional ML-based WF attacks [4, 7, 9]. With zero latency overhead and less than 60% bandwidth overhead, WTF-PAD offers a more practical solution than previous defenses. Cui et al. [54] introduced an algorithm that generates realistic cover traffic to mitigate WF attacks. The proposed method uses a user’s historical network traffic to generate cover traffic that mimics their browsing habits, making it difficult for adversaries to distinguish real traffic from noise. Cui et al. [164] introduced an improved algorithm that also uses historical browsing data to generate realistic cover traffic, reducing attack accuracy to 14%–16% in both simulations and real-world experiments. Gong et al. [57] presented two lightweight defenses, FRONT and GLUE, which aim to protect against advanced WF attacks [9, 11, 93, 36] with minimal cost. FRONT obfuscates the beginning of traffic traces by injecting randomized dummy packets, while GLUE inserts dummy packets between traces to create the illusion of a single, continuous trace. Both methods operate without delaying client packets and require only a small number of dummy packets, offering a practical and efficient defense.
@@
Pulls [55] introduced Interspace, a promising padding mechanism for Tor that offers a tunable balance between defense strength and bandwidth overhead in 2020. This work emphasizes the importance of probabilistic defenses and highlights the need for further research to overcome real-world deployment challenges, such as timing inaccuracies and adaptive attackers. Abusnaina et al. [67] proposed the Deep Fingerprinting Defender (DFD), a defense mechanism targeting deep learning-based WF attacks. DFD employs per-burst dummy packet injection in both one-way and two-way modes, disrupting traffic patterns inherent in Tor user traces. Luo et al. [160] introduced Random Bidirectional Padding (RBP) that combines time sampling with bidirectional dummy packet injection to obfuscate traffic features. RBP effectively disrupts inter-arrival time (IAT) features and packet distributions, making it more difficult for classifiers to extract meaningful patterns.
@@
The protocol layer metadata (e.g., cell sequences and timing) of Tor enables accurate classification of circuit types, such as onion service circuits versus regular clearnet circuits. To address this issue, Kadianakis et al. [68] presented the first comprehensive framework and adaptive padding strategies for defending against circuit fingerprinting in Tor. By exploring both delay-based and preemptive padding defenses, the authors demonstrate that practical, low-cost defenses can thwart powerful circuit fingerprinting adversaries. The work also opens new directions for integrating padding into Tor to protect against broader classes of traffic analysis attacks.
@@
To reduce the requirement of high computation or additional infrastructure for previous WF defense works, Hong et al. [78] introduced Randomized Delaying and Padding (RanDePad), a WF defense mechanism designed to resist WF attacks while maintaining low latency and controllable bandwidth. RanDePad uses adaptive random delaying and adaptive random padding to disrupt the time distribution patterns of traffic. The key contribution is its ability to achieve a superior balance of low latency and controllable bandwidth compared to other state-of-the-art WF defenses, while effectively reducing the accuracy of DL-based attacks. Wang et al. [162] introduced Tree-structured Burst-sequence Padding (TBP), a defense mechanism that inserts dummy burst sequences to mimic fake resources. The method uses multiple random distributions to randomize both the timing and direction of traffic bursts, thereby obfuscating the user’s online activity. This approach obscures intra-class similarities and inter-class differences, reducing the true positive rate (TPR) of advanced CNN-based attacks such as Var-CNN to 44.7% with a 55% bandwidth overhead, outperforming defenses like FRONT by 15% in TPR reduction. Lu et al. [72] introduced WF Defense based on Distribution Distance Padding (WFD3P), a lightweight, zero-delay defense mechanism designed to combat WF attacks. WFD3P selects target traces from a pool with the maximum distributional distance from the user’s traffic and injects dummy packets based on the selected trace’s characteristics. Unlike global-pattern padding methods, WFD3P dynamically tailors padding to each source trace, improving defense granularity and effectiveness.
@@
In 2023, Yang et al. [161] proposed Advanced Adaptive Padding (AAP), a zero-delay, moderate-overhead defense that extends the original Adaptive Padding (AP) approach. The AAP focuses on obfuscating HTTP burst patterns during website loading, providing enhanced protection for the Tor anonymity network. McGuan et al. [163] introduced Beta Randomized Obfuscation (BRO), a lightweight and zero-delay defense against WF attacks. BRO uses a beta distribution to strategically inject dummy packets, especially targeting the feature-rich beginning of traffic traces. With minimal bandwidth overhead (16.8–25.4%), BRO enhances trace randomness and dummy packet clustering, reducing the accuracy of state-of-the-art attacks like Deep Fingerprinting (DF) to 43%, compared to 54.2% under FRONT. Pulls et al. [165] introduced Maybenot, a flexible framework designed to implement traffic analysis defenses against metadata-based WF attacks in encrypted communication protocols like Tor, VPNs, and QUIC. Maybenot aims to provide a generalized, lightweight, and adaptable solution for deploying defenses without requiring extensive modifications to existing protocols. Huang and Du [166] proposed Break-Pad, which injects padding packets when the number of consecutive incoming packets exceeds a threshold, thereby breaking long bursts into smaller ones. In addition, Break-Pad uses randomized thresholds and padding counts sampled from probability distributions (Weibull for thresholds, Pareto for padding counts) to enhance unpredictability.
@@ V-B2 Regularization-based defenses
Even though adaptive padding-based mechanisms can efficiently defend against WF attacks by adding packet padding or dummy packets, attackers can still accurately estimate the accessed website using DL-based attack models through interval timing analysis and deep learning-based feature extraction. Because the padding scheme is often deterministic or follows predictable rules [107]. To defend against DL-based WF attacks effectively, regularization-based defenses intentionally reduce model performance on WF attacks by regulating the data representation, thereby diminishing the model’s ability to learn discriminative features. As shown in Figure 8, the regularization-based defense approach enforces fixed rules and patterns, such as uniform inter-packet intervals or consistent burst sizes, that all web traffic must follow [167]. TABLE XI summarizes representative regularization-based defenses.
@@
In 2012, Dyer et al. [58] conducted the first comprehensive analysis of low-level countermeasures, such as per-packet padding, in the context of closed-world WF attacks. Their work introduced Buffered Fixed-Length Obfuscation (BuFLO), a defense that sends fixed-size packets at fixed intervals by adding dummy traffic, thereby obfuscating total traffic duration and bandwidth. BuFLO was shown to reduce WF attack [8, 66, 4, 105] accuracy by up to 5%, but its overhead was considered impractically high for real-world deployment. To address the inefficiency of BuFLO, Cai et al. [69] introduced Congestion-Sensitive BuFLO (CS-BuFLO) that adds rate adaptation, congestion sensitivity, and practical padding control. It offers strong protection with real-world deployability, though it still incurs higher overhead than lightweight defenses. In the same year, Cai et al. [15] introduced another defense method, Tamaraw, to achieve a better security/bandwidth trade-off than BuFLO. This method, combined with a theoretical framework and feature-based testing, offers a blueprint for building future defenses that are efficient and attack-agnostic. But the Tamaraw can only achieve non-uniform accuracy.
@@
Later, Lu et al. [56] introduced DynaFlow, a WF defense mechanism by dynamically adjusting website traffic flows using the average inter-packet arrival time from the previous period. DynaFlow enhances security while improving efficiency by over 40% compared to previous WF defense mechanisms [15, 58, 69], eliminating the need for a pre-established database and extending protection to dynamically-generated websites. Experimental evaluations demonstrate that DynaFlow significantly reduces attacker accuracy while maintaining acceptable latency and bandwidth overheads, making it a promising and practical solution for safeguarding user privacy amid increasing surveillance. To address the challenges of high bandwidth and latency overhead, as well as the reliance on additional infrastructure, Holland et al. [73] proposed RegularTor. RegularTor mitigates these attacks by leveraging common patterns in typical web browsing traffic, thereby reducing both the resource overhead of the defense and the accuracy of state-of-the-art WF classifiers. Experimental results demonstrate that RegulaTor reduces the accuracy of the Tik-Tok attack from 66% to 25.4%, significantly outperforming prior defense methods. To effectively defend against deep learning-based WF attacks with lower overhead, Chen et al. [168] proposed Timing-Assistant Traffic Burst Shaping (BM). This approach addresses the limitations of existing defenses, such as WTF-PAD and Walkie-Talkie, by incorporating temporal features. BM redefines traffic bursts using dynamic time thresholds and merges sequences from sensitive and non-sensitive websites into a unified “super-sequence” to obscure burst-level patterns.
@@ V-B3 Traffic Morphing-Based Defenses
Website morphing is another privacy-enhancing technique designed to defend against WF attacks. This mechanism will transform the observable traffic characteristics (e.g., packet size distributions) of a webpage (the source) so it resembles another webpage (the target), thus misleading classifiers without excessive overhead [105, 17]. By adjusting specific traffic features, website morphing effectively shifts the classifier’s prediction away from the true class and toward a predefined, misleading target class. Figure 9 illustrates the traffic morphing mechanism, which adds padding or dummy cells to both incoming and outgoing traffic, transforming the packet patterns to resemble those of a different website. TABLE XII shows a comprehensive comparison of the representative traffic morphing-based defenses.
@@
In 2014, Supersequence [9] was introduced to transform the actual packet sequence of a website into a common supersequence shared across an anonymity set (e.g., a group of packet sequences from different web pages). To implement this, the authors approximate the Shortest Common Supersequence (SCS) across packet sequences in an anonymity set. They cluster packet traces into sets that share similar characteristics and generate a supersequence for each set. When a page is visited, its trace is padded or reshaped to match the supersequence, preventing attackers from identifying the specific site visited. To address the problem of hampering the user experience and burdening the server with large overheads for previous works, Wang and Goldberg [45] presented Walkie-Talkie, which is designed to make sensitive and non-sensitive pages look the same. This approach employs half-duplex communication to group outgoing and incoming packets, along with burst molding to add fake packets, making it difficult for attackers to distinguish between sensitive and non-sensitive web page accesses. The authors demonstrate that Walkie-Talkie [45] achieves lower bandwidth and time overhead than existing defenses, effectively mitigating multi-page attacks. However, Walkie-Talkie requires prior knowledge of each site and is less effective for web pages with dynamic content.
@@
Chan-Tin et al. [22] introduced a mechanism against WF attacks by leveraging traffic morphing and clustering to make websites within the same cluster indistinguishable. The proposed method groups websites based on their network traffic patterns (packet sizes and counts) and morphs all traffic in a cluster to resemble the largest website in that cluster, significantly reducing the attacker’s ability to identify specific sites. Al-Naami et al. [23] introduced BiMorphing, a technique that utilizes bi-directional burst sampling and optimization methods to obfuscate traffic patterns. This approach achieves minimal bandwidth overhead and zero-delay transmission. The defense is evaluated to outperform previous methods in both closed-world and open-world scenarios. Previous defenses [69, 56, 45, 23] often struggle to counter DL-based WF attacks effectively without introducing substantial bandwidth or latency overhead. To solve this issue, Ling et al. [24] presented WFGuard, a traffic morphing defense technique against DL-based WF attacks based on fuzzing testing. WFGuard leverages fine-grained neuron information from WF classifiers to design a joint optimization function, which is then optimized using gradient ascent to simultaneously maximize both neuron values and the likelihood of misclassification in DL-based WF classifiers. During each traffic mutation cycle, a gradient-based approach is employed to generate dummy traffic injection patterns, continuously evolving the traffic until a pattern is found that successfully deceives the classifier. Extensive evaluations show that WFGuard significantly reduces the accuracy of DL-based WF classifiers (e.g., DF and Var-CNN) to as low as 4.43%, while only incurring a modest 11.04% bandwidth overhead. In the same year, Shen et al. [19] presented Palette, a practical defense against WF attacks that threaten user privacy in Tor. Unlike traditional defenses, Palette focuses on traffic cluster anonymization by grouping websites with similar traffic patterns into clusters and regulating their traffic to conform to a uniform pattern. This approach makes it significantly harder for attackers to distinguish between websites within the same cluster, effectively mitigating the risk posed by modern deep learning-based WF attacks.

### v2

@@ V-B Survey on WF Defenses
Tor encrypts and routes traffic through a network of relays to obscure user identity and destination, but it does not conceal traffic bursts or timing information, leaving it vulnerable to WF attacks. These attacks are particularly concerning in Tor, where users rely on strong privacy guarantees [giovanni2017bayes]. WF attacks against the first-hop (entry) relay can accurately infer user activity, especially in closed-world settings where the set of possible websites is predefined. To address this threat, researchers have proposed various WF defenses aimed at reducing the distinguishability of traffic patterns across websites, thereby diminishing the effectiveness of WF classifiers.
@@ V-B1 Adaptive Padding-Based Defenses
Adaptive traffic padding is a technique used to protect against WF attacks by dynamically injecting dummy traffic and introducing controlled delays into network traffic streams [shmatikov2006timing]. This dynamic behavior can be achieved by traffic obfuscation or randomized padding. Traffic obfuscation involves inserting dummy packets into statistically improbable timing gaps or low-activity regions of the traffic stream, making it more difficult for an attacker to identify meaningful patterns. Randomized padding further enhances this effect by varying the timing and size of the inserted packets, thereby increasing uncertainty for WF classifiers. This adaptive approach enables the system to disrupt fingerprinting attacks while maintaining low overhead in terms of bandwidth and latency [witwer2022padding]. Figure 7 illustrates the mechanism of adaptive padding. In the figure, coral rectangles represent outgoing packets, while light blue rectangles denote incoming packets. The light yellow and light green rectangles indicate the dummy packets added for outgoing and incoming traffic, respectively. The insertion of these dummy packets helps obscure burst patterns, timing information, and other features commonly exploited by WF attacks.
@@
TABLE X summarizes existing representative adaptive padding-based defenses, highlighting their strengths and limitations to facilitate comparison for researchers. Juarez et al. [juarez2016toward] introduced WTF-PAD, an adaptive padding defense that protects Tor users from WF attacks. This defense employs adaptive padding to disrupt distinctive traffic patterns, thereby significantly reducing the accuracy of traditional ML-based WF attacks [panchenko2011website, cai2012touching, wang2014effective]. With zero latency overhead and less than 60% bandwidth overhead, WTF-PAD offers a more practical solution than previous defenses. Cui et al. [cui2018realistic] introduced an algorithm that generates realistic cover traffic to mitigate WF attacks. The proposed method uses a user’s historical network traffic to generate cover traffic that mimics their browsing habits, making it difficult for adversaries to distinguish real traffic from noise. Cui et al. [cui2019efficient] introduced an improved algorithm that also uses historical browsing data to generate realistic cover traffic, reducing attack accuracy to 14%–16% in both simulations and real-world experiments. Gong et al. [gong2020zero] presented two lightweight defenses, FRONT and GLUE, which aim to protect against advanced WF attacks [wang2014effective, panchenko2016website, sirinam2018deep, hayes2016k] with minimal cost. FRONT obfuscates the beginning of traffic traces by injecting randomized dummy packets, while GLUE inserts dummy packets between traces to create the illusion of a single, continuous trace. Both methods operate without delaying client packets and require only a small number of dummy packets, offering a practical and efficient defense.
@@
Pulls [pulls2020towards] introduced Interspace, a promising padding mechanism for Tor that offers a tunable balance between defense strength and bandwidth overhead in 2020. This work emphasizes the importance of probabilistic defenses and highlights the need for further research to overcome real-world deployment challenges, such as timing inaccuracies and adaptive attackers. Abusnaina et al. [abusnaina2020dfd] proposed the Deep Fingerprinting Defender (DFD), a defense mechanism targeting deep learning-based WF attacks. DFD employs per-burst dummy packet injection in both one-way and two-way modes, disrupting traffic patterns inherent in Tor user traces. Luo et al. [luo2021rbp] introduced Random Bidirectional Padding (RBP) that combines time sampling with bidirectional dummy packet injection to obfuscate traffic features. RBP effectively disrupts inter-arrival time (IAT) features and packet distributions, making it more difficult for classifiers to extract meaningful patterns.
@@
The protocol layer metadata (e.g., cell sequences and timing) of Tor enables accurate classification of circuit types, such as onion service circuits versus regular clearnet circuits. To address this issue, Kadianakis et al. [kadianakis2021tor] presented the first comprehensive framework and adaptive padding strategies for defending against circuit fingerprinting in Tor. By exploring both delay-based and preemptive padding defenses, the authors demonstrate that practical, low-cost defenses can thwart powerful circuit fingerprinting adversaries. The work also opens new directions for integrating padding into Tor to protect against broader classes of traffic analysis attacks.
@@
To reduce the requirement of high computation or additional infrastructure for previous WF defense works, Hong et al. [hong2022website] introduced Randomized Delaying and Padding (RanDePad), a WF defense mechanism designed to resist WF attacks while maintaining low latency and controllable bandwidth. RanDePad uses adaptive random delaying and adaptive random padding to disrupt the time distribution patterns of traffic. The key contribution is its ability to achieve a superior balance of low latency and controllable bandwidth compared to other state-of-the-art WF defenses, while effectively reducing the accuracy of DL-based attacks. Wang et al. [wang2023tbp] introduced Tree-structured Burst-sequence Padding (TBP), a defense mechanism that inserts dummy burst sequences to mimic fake resources. The method uses multiple random distributions to randomize both the timing and direction of traffic bursts, thereby obfuscating the user’s online activity. This approach obscures intra-class similarities and inter-class differences, reducing the true positive rate (TPR) of advanced CNN-based attacks such as Var-CNN to 44.7% with a 55% bandwidth overhead, outperforming defenses like FRONT by 15% in TPR reduction. Lu et al. [lu2023lightweight] introduced WF Defense based on Distribution Distance Padding (WFD3P), a lightweight, zero-delay defense mechanism designed to combat WF attacks. WFD3P selects target traces from a pool with the maximum distributional distance from the user’s traffic and injects dummy packets based on the selected trace’s characteristics. Unlike global-pattern padding methods, WFD3P dynamically tailors padding to each source trace, improving defense granularity and effectiveness.
@@
In 2023, Yang et al. [yang2023aap] proposed Advanced Adaptive Padding (AAP), a zero-delay, moderate-overhead defense that extends the original Adaptive Padding (AP) approach. The AAP focuses on obfuscating HTTP burst patterns during website loading, providing enhanced protection for the Tor anonymity network. McGuan et al. [mcguan2024practical] introduced Beta Randomized Obfuscation (BRO), a lightweight and zero-delay defense against WF attacks. BRO uses a beta distribution to strategically inject dummy packets, especially targeting the feature-rich beginning of traffic traces. With minimal bandwidth overhead (16.8–25.4%), BRO enhances trace randomness and dummy packet clustering, reducing the accuracy of state-of-the-art attacks like Deep Fingerprinting (DF) to 43%, compared to 54.2% under FRONT. Pulls et al. [pulls2023maybenot] introduced Maybenot, a flexible framework designed to implement traffic analysis defenses against metadata-based WF attacks in encrypted communication protocols like Tor, VPNs, and QUIC. Maybenot aims to provide a generalized, lightweight, and adaptable solution for deploying defenses without requiring extensive modifications to existing protocols. Huang and Du [huang2024break] proposed Break-Pad, which injects padding packets when the number of consecutive incoming packets exceeds a threshold, thereby breaking long bursts into smaller ones. In addition, Break-Pad uses randomized thresholds and padding counts sampled from probability distributions (Weibull for thresholds, Pareto for padding counts) to enhance unpredictability.
@@ V-B2 Regularization-based defenses
Even though adaptive padding-based mechanisms can efficiently defend against WF attacks by adding packet padding or dummy packets, attackers can still accurately estimate the accessed website using DL-based attack models through interval timing analysis and deep learning-based feature extraction. Because the padding scheme is often deterministic or follows predictable rules [wang2021one]. To defend against DL-based WF attacks effectively, regularization-based defenses intentionally reduce model performance on WF attacks by regulating the data representation, thereby diminishing the model’s ability to learn discriminative features. As shown in Figure 8, the regularization-based defense approach enforces fixed rules and patterns, such as uniform inter-packet intervals or consistent burst sizes, that all web traffic must follow [yang2024website]. TABLE XI summarizes representative regularization-based defenses.
@@
In 2012, Dyer et al. [dyer2012peek] conducted the first comprehensive analysis of low-level countermeasures, such as per-packet padding, in the context of closed-world WF attacks. Their work introduced Buffered Fixed-Length Obfuscation (BuFLO), a defense that sends fixed-size packets at fixed intervals by adding dummy traffic, thereby obfuscating total traffic duration and bandwidth. BuFLO was shown to reduce WF attack [herrmann2009website, liberatore2006inferring, panchenko2011website, wright2009traffic] accuracy by up to 5%, but its overhead was considered impractically high for real-world deployment. To address the inefficiency of BuFLO, Cai et al. [cai2014cs] introduced Congestion-Sensitive BuFLO (CS-BuFLO) that adds rate adaptation, congestion sensitivity, and practical padding control. It offers strong protection with real-world deployability, though it still incurs higher overhead than lightweight defenses. In the same year, Cai et al. [cai2014systematic] introduced another defense method, Tamaraw, to achieve a better security/bandwidth trade-off than BuFLO. This method, combined with a theoretical framework and feature-based testing, offers a blueprint for building future defenses that are efficient and attack-agnostic. But the Tamaraw can only achieve non-uniform accuracy.
@@
Later, Lu et al. [lu2018dynaflow] introduced DynaFlow, a WF defense mechanism by dynamically adjusting website traffic flows using the average inter-packet arrival time from the previous period. DynaFlow enhances security while improving efficiency by over 40% compared to previous WF defense mechanisms [cai2014systematic, dyer2012peek, cai2014cs], eliminating the need for a pre-established database and extending protection to dynamically-generated websites. Experimental evaluations demonstrate that DynaFlow significantly reduces attacker accuracy while maintaining acceptable latency and bandwidth overheads, making it a promising and practical solution for safeguarding user privacy amid increasing surveillance. To address the challenges of high bandwidth and latency overhead, as well as the reliance on additional infrastructure, Holland et al. [holland2020regulator] proposed RegularTor. RegularTor mitigates these attacks by leveraging common patterns in typical web browsing traffic, thereby reducing both the resource overhead of the defense and the accuracy of state-of-the-art WF classifiers. Experimental results demonstrate that RegulaTor reduces the accuracy of the Tik-Tok attack from 66% to 25.4%, significantly outperforming prior defense methods. To effectively defend against deep learning-based WF attacks with lower overhead, Chen et al. [chen2023web] proposed Timing-Assistant Traffic Burst Shaping (BM). This approach addresses the limitations of existing defenses, such as WTF-PAD and Walkie-Talkie, by incorporating temporal features. BM redefines traffic bursts using dynamic time thresholds and merges sequences from sensitive and non-sensitive websites into a unified “super-sequence” to obscure burst-level patterns. Shen et al. [shen2024real] presented Palette, a practical defense against WF attacks that threaten user privacy in Tor. Unlike traditional defenses, Palette focuses on traffic cluster anonymization by grouping websites with similar traffic patterns into clusters and regulating their traffic to conform to a uniform pattern. This approach makes it significantly harder for attackers to distinguish between websites within the same cluster, effectively mitigating the risk posed by modern deep learning-based WF attacks.
@@ V-B3 Traffic Morphing-Based Defenses
Website morphing is another privacy-enhancing technique designed to defend against WF attacks. This mechanism will transform the observable traffic characteristics (e.g., packet size distributions) of a webpage (the source) so it resembles another webpage (the target), thus misleading classifiers without excessive overhead [wright2009traffic, nithyanand2014glove]. By adjusting specific traffic features, website morphing effectively shifts the classifier’s prediction away from the true class and toward a predefined, misleading target class. Figure 9 illustrates the traffic morphing mechanism, which adds padding or dummy cells to both incoming and outgoing traffic, transforming the packet patterns to resemble those of a different website. TABLE XII shows a comprehensive comparison of the representative traffic morphing-based defenses.
@@
In 2014, Supersequence [wang2014effective] was introduced to transform the actual packet sequence of a website into a common supersequence shared across an anonymity set (e.g., a group of packet sequences from different web pages). To implement this, the authors approximate the Shortest Common Supersequence (SCS) across packet sequences in an anonymity set. They cluster packet traces into sets that share similar characteristics and generate a supersequence for each set. When a page is visited, its trace is padded or reshaped to match the supersequence, preventing attackers from identifying the specific site visited. To address the problem of hampering the user experience and burdening the server with large overheads for previous works, Wang and Goldberg [wang2017walkie] presented Walkie-Talkie, which is designed to make sensitive and non-sensitive pages look the same. This approach employs half-duplex communication to group outgoing and incoming packets, along with burst molding to add fake packets, making it difficult for attackers to distinguish between sensitive and non-sensitive web page accesses. The authors demonstrate that Walkie-Talkie [wang2017walkie] achieves lower bandwidth and time overhead than existing defenses, effectively mitigating multi-page attacks. However, Walkie-Talkie requires prior knowledge of each site and is less effective for web pages with dynamic content.
@@
Chan-Tin et al. [chan2018website] introduced a mechanism against WF attacks by leveraging traffic morphing and clustering to make websites within the same cluster indistinguishable. The proposed method groups websites based on their network traffic patterns (packet sizes and counts) and morphs all traffic in a cluster to resemble the largest website in that cluster, significantly reducing the attacker’s ability to identify specific sites. Al-Naami et al. [al2019bimorphing] introduced BiMorphing, a technique that utilizes bi-directional burst sampling and optimization methods to obfuscate traffic patterns. This approach achieves minimal bandwidth overhead and zero-delay transmission. The defense is evaluated to outperform previous methods in both closed-world and open-world scenarios. Previous defenses [cai2014cs, lu2018dynaflow, wang2017walkie, al2019bimorphing] often struggle to counter DL-based WF attacks effectively without introducing substantial bandwidth or latency overhead. To solve this issue, Ling et al. [ling2024wfguard] presented WFGuard, a traffic morphing defense technique against DL-based WF attacks based on fuzzing testing. WFGuard leverages fine-grained neuron information from WF classifiers to design a joint optimization function, which is then optimized using gradient ascent to simultaneously maximize both neuron values and the likelihood of misclassification in DL-based WF classifiers. During each traffic mutation cycle, a gradient-based approach is employed to generate dummy traffic injection patterns, continuously evolving the traffic until a pattern is found that successfully deceives the classifier. Extensive evaluations show that WFGuard significantly reduces the accuracy of DL-based WF classifiers (e.g., DF and Var-CNN) to as low as 4.43%, while only incurring a modest 11.04% bandwidth overhead.